# Improving Video Generation with Human Feedback

### v1

@@ 4.3 Noisy Reward Guidance
VideoScore (He et al., 2024) 49.03 71.69 41.80 50.22 47.41 47.72 59.05 51.09 37.24 50.34
LiFT* (Wang et al., 2024b) 37.06 58.39 39.08 57.26 47.53 55.97 59.04 54.91 33.79 55.43
VisionRewrd (Xu et al., 2024a) 51.56 72.41 56.77 67.59 47.43 59.03 59.03 60.98 46.56 61.15
@@
Table 3: Ablation study on reward model type, seprate tokens and data augmentation. Bold: Best Performance.
Variants RM Type Separate Tokens GenAI-Bench VideoGen-RewardBench
II BT 47.74 71.21 61.22 72.58 52.33 77.10 59.43 73.50 53.06 71.62
III BTT 48.27 70.89 61.50 73.39 60.52 76.31 64.64 72.40 53.55 72.12
IV BTT 49.41 72.89 61.26 73.59 59.68 75.66 66.03 74.70 53.80 72.20
where the expectation is taken over $({\bm{x}}^{w},{\bm{x}}^{l},{\bm{y}})\sim\mathcal{D},\,t\sim U(0,1)$. In this way, the reward function $r_{\theta}(\cdot,t)$ learns to preserve the relative preference between noised videos at any noise level $t$. However, advanced video generation models typically work in a latent space and rely on a VAE (Kingma, 2013) to decode the latent representation back into a video. In Eq. 15, computing the gradient of the reward with respect to ${\bm{x}}_{t}$ would require first decoding the latent representation into a video and then passing the gradient through the large VAE decoder back to the latent space, incurring high computational cost and memory usage.
To address this challenge, we propose training the reward model directly within the latent space using a pretrained VDM. Since the VDM has already been trained on noisy latent representations, we can develop a lightweight, time-dependent reward function by utilizing only the initial layers of the pretrained model rather than the entire VDM. Subsequently, we apply Equation 15 at each inference step to implement reward-guided generation.
5 Experiments
In this section, we empirically evaluate our reward model and reinforcement learning alignment algorithms on Text-to-Video (T2V) tasks.
@@ 5.1 Reward Learning
We utilize Qwen2-VL-2B (Wang et al., 2024a) as the backbone for our reward model and train it with BTT loss. To finetune the model, LoRA (Hu et al., 2021) is applied to update all linear layers in the language model, while the vision encoder’s parameters are fully optimized. The training process is conducted with a batch of 32 and a learning rate of $2\times 10^{-6}$, with the model trained over two epochs. This setup requires approximately 72 A800 GPU hours. Several observations were made during training. First, higher video resolution and more frames generally improved the reward model’s performance. Second, using a stable sampling interval instead of a fixed frame number significantly enhanced the accuracy of motion quality evaluations, particularly when trained with videos of varying durations. In practice, we sample videos at 2 fps, with a resolution of approximately $448\times 448$ pixels during the training process. The original aspect ratio is preserved to ensure visual quality.
@@
To assess the performance of our reward model in evaluating T2V generation, we employ a comprehensive evaluation framework designed to ensure a fair comparison across different models and mitigate potential domain biases. We utilize two primary benchmarks for evaluation. (1). VideoGen-RewardBench: Derived from the third-party prompt-video dataset VideoGen-Eval (Zeng et al., 2024), VideoGen-RewardBench is specifically designed to evaluate modern T2V models. To address the absence of human annotations in VideoGen-Eval, we manually construct 26.5k triplets and hire annotators to provide pairwise preference labels, following the procedure outlined in Sec. 3.1. Additionally, annotators assess the overall quality between the two videos, serving as a universal label for reward models across various dimensions. (2). GenAI-Bench (Jiang et al., 2024b): : Serving as a complementary benchmark, GenAI-Bench features short videos (2 seconds in duration) generated by pre-Sora-era T2V models. This benchmark enables us to evaluate our reward model’s capability in assessing earlier-generation models, despite their limited representation in our training set. We summarize the key difference between VideoGen-RewardBench and GenAI-Bench in Appendix D.1.
We compare our reward model against existing baselines, including VideoScore (He et al., 2024), as well as two concurrent works: LiFT (Wang et al., 2024b), and VisionReward (Xu et al., 2024a). Consistent with prevailing practices in LLMs (Lambert et al., 2024), we employ pairwise accuracy as a primary metric. Inspired by VisionReward (Xu et al., 2024a), we report two accuracy metrics: ties-included accuracy (Deutsch et al., 2023) and ties-excluded accuracy. We calculate the overall accuracy on GenAI-Bench and both overall and dimension-specific (VQ, MQ, TA) accuracy on VideoGen-RewardBench. Additional details of the evaluation across different methods are provided in the Appendix D.2.
@@
Main Results.
Tab 2 presents the pairwise accuracy of various reward models on GenAI-Bench (representing pre-Sora-era models) and VideoGen-RewardBench (representing modern T2V models). VideoScore is primarily designed to evaluate videos from earlier-generation models. It performs well on GenAI-Bench but exhibits close-to-random predictions on VideoGen-RewardBench, reflecting its inability to generalize to the improved capabilities of modern T2V models. LiFT adopts the VLM-as-a-judge approach to train its reward model, but it struggles with pairwise differentiation. This is because LiFT tends to assign similar scores to both videos in a pair. When calculating its ties-excluded accuracy, we randomly convert ties prediction to chosen/rejected categories with a 0.5 probability, indicating the model fail to distinguish between video qualities. VisionReward demonstrates competitive performance on GenAI-Bench but shows diminished performance on VideoGen-RewardBench, particularly in ties-included accuracy for visual quality and motion quality. This shortfall can be attributed to the significant improvements in visual fidelity and motion smoothness exhibited by modern T2V models, which VisionReward struggles to effectively assess. In contrast, our method VideoReward outperforms all other models on VideoGen-RewardBench, showcasing its effectiveness in evaluating outputs from state-of-the-art T2V models. Moreover, despite being trained on a distinct dataset, our approach also achieves comparable performance to VisionReward on GenAI-Bench, highlighting its strong generalization across different eras of T2V models.
Ablation Study.
As discussed in Sec. 3.2, we have highlighted several key design choices for our reward model and demonstrated their effectiveness on the validation set. To further prove these findings, we conduct an ablation study on the evaluation benchmark, providing a quantitative supplement to our analysis. Results are demonstrates in Tab. 3.
Regarding the type of reward model, we compare three variants: the regression-style reward model, the Bradley-Terry reward model, and the Bradley-Terry-With-Tied reward model. The BT reward model shows a slight advantage over the regression-style reward model, likely due to inherent strengths of pairwise annotations. Pairwise annotations more accurately capture the ranking relationships within the same context and demonstrate greater tolerance to potential noise in data annotations. The BTT reward model achieves comparable performance to the BT model on ties-excluded accuracy, but exhibits a significant improvement in ties-included accuracy. This improvement stems from the BTT model’s explicit modeling to tied pairs, enabling it to learn a more robust decision boundary that effectively captures neutral relationships in ambiguous cases.
Beyond the benefits of disentangling reward’s attribute as discussed in Sec. 3.2, we observe that adopting separate tokens rather than a shared last token further enhances the model’s overall performance. This design allows the model to better represent distinct aspects of the reward, thus improving alignment with human preferences.
Figure 5: Visual comparison of videos generated by the original pretrained model and the Flow-DPO aligned model.
5.2 Video Alignment
Training Setting.
Our pretrained model $p_{\text{ref}}$ is a transformer-based latent flow model trained using rectified flow. In all alignment experiments, we applied LoRA to fine-tune the transformer models’ linear layers, as our findings indicate that full parameter fine-tuning can degrade the model’s performance or potentially lead to model collapse. For supervised fine-tuning (SFT), we utilize only the “chosen data.” In the RWR experiments, we first calculate the mean and variance of the rewards across all training samples and normalize the rewards to have a mean of 0 and a variance of 1. For the VDM reward model, we extract and concatenate the features from the first 20 transformer blocks of our VDM. These concatenated features are then fed into three fully connected heads, which output a three-dimensional reward. All parameters of the first 20 transformer blocks are trained. For the training dataset, following the approach of Rafailov et al. (Rafailov et al., 2024), we employ VideoReward as the ground-truth reward model to simulate human feedback and relabel our training dataset. Consequently, T2V models trained on this synthetically relabeled dataset can be reliably evaluated using VideoReward , ensuring a fair assessment of their performance. More experimental details, including LoRA architectures, VDM reward model and final hyperparameter settings, are provided in the Appendix F.
Figure 6: Human evaluation of Flow-DPO aligned model vs. pretrained model on VideoGen-Eval, which contains 400 prompts.
Evaluation.
We evaluate performance using both automatic and human assessments. The automatic evaluation comprises the win rate, as determined by VideoReward , and the Vbench score. The win rate is calculated by having VideoReward assign scores to outputs from both the pretrained and aligned models, then determining the proportion of instances where the aligned model’s reward exceeds that of the pretrained model. Since VideoReward generates continuous floating-point scores, ties do not occur. Vbench serves as a fine-grained benchmark for evaluating T2V models from two main perspectives: Quality and Semantic. The Quality aspect focuses on the perceptual quality of the synthesized video without considering the input condition, corresponding to our VQ and MQ metrics. The Semantic aspect assesses whether the synthesized video aligns with the user-provided guiding condition, corresponding to TA. For human evaluation, each sample is independently reviewed by two annotators to ensure reliability. In cases of disagreement, a third annotator is consulted to resolve conflicts. We employ same random seeds to generate videos for each method, ensuring a fair comparison. We employ Vbench, VideoGen-Eval and TA-Hard prompts for evaluation. We find that the standard Vbench and VideoGen-Eval prompts are relatively straightforward in terms of text alignment, as the pretrained model consistently generates well-aligned video content. To create a more rigorous assessment, we have developed additional challenging prompts (TA-Hard) that feature combinations of two subjects and incorporate uncommon actions. For example: “A talking apple with eyes and a mouth, and a singing banana with legs hosting a talent show in a vibrant theater.” We provide a prompt subset of TA-Hard in Appendix H.
Multi-dimensional Alignment.
We employ linear scalarization (Li et al., 2020), defined as $r=\frac{1}{3}(r_{\text{vq}}+r_{\text{mq}}+r_{\text{ta}})$, to consolidate multi-dimensional preferences into a single score. For each prompt ${\bm{y}}$, the reward model annotates preferences between two generated videos, such that ${\bm{x}}^{w}$ is preferred over ${\bm{x}}^{l}$ if $r({\bm{x}}^{w},{\bm{y}})>r({\bm{x}}^{l},{\bm{y}})$. This process generates a preference dataset $D=\{({\bm{y}},{\bm{x}}^{w},{\bm{x}}^{l})\}$. As presented in Tab. 4, Flow-DPO with a constant $\beta$ achieves significant improvements over the pretrained model across three dimensions and outperforms both SFT and Flow-RWR. Conversely, Flow-DPO with a timestep-dependent $\beta$ exhibits reduced performance in the TA task compared to the pretrained model. Although it maintains high reward win rates in VQ and MQ, we observe severe reward hacking issues in the generated videos. In contrast, Flow-DPO with a constant $\beta$ also achieves high VQ and MQ scores without encountering reward hacking problems. This suggests that Flow-DPO with a timestep-dependent $\beta$ struggles to effectively learn the TA task compared to VQ and MQ tasks. We will discuss this issue in detail in Section 5.25.2. Fig. 610 demonstrates that the aligned model using Flow-DPO outperforms the pretrained model in human evaluations for both VideoGen-Eval and TA-Hard prompts.
@@ 5.2 Video Alignment
We also investigate the ability of different methods to align the T2V model on a specific task: TA. As shown in Tab. 5, Flow-DPO with a constant $\beta$ achieves best performance across all datasets.
Table 6: Reward guidance using multi-dimensional rewards on VideoGen-Eval prompts. The weighted reward guidence allows users to apply arbitrary weightings to multiple alignment objectives during inference to meet personalized user requirements.
@@
Reward Guidance.
We also employ linear scalarization (Li et al., 2020) to compute a weighted sum of rewards for each dimension and backpropagate the gradients to the noised latent space. Tab. 6 demonstrates that reward guidance can improve the pretrained model in the corresponding dimensions when using the respective dimension’s reward to guide the generation process. Furthermore, the weighted reward approach simultaneously enhances performance across all dimensions.
Figure 7: Accuracy of time-dependent $\beta_{t}$ vs. constant $\beta$ for TA: Flow-DPO with a constant $\beta$ consistently outperforms the timestep-dependent $\beta$ across various settings.
@@
We meticulously adapted diffusion-DPO to flow-based models, resulting in Equation 4.1, where $\beta_{t}=\beta(1-t)^{2}$. As illustrated in Fig. 7, under various $\beta$ settings, Flow-DPO with a constant $\beta$ consistently outperforms Flow-DPO with a timestep dependent $\beta$. We note that none of the results in the figure exhibit reward hacking issues. Timestep dependent $\beta$ may cause uneven training across different timesteps (Hang et al., 2023), since T2V models utilize shared model weights for various noise levels.
Ablation on Adding Noise to the VDM Latent Reward Model.
Tab. 7 presents the results of applying reward guidance exclusively with MQ rewards on TA-Hard prompts. The reward model trained with noised latents effectively guides the generation process, whereas the reward model trained solely on cleaned latents fails to provide meaningful gradient guidance for intermediate noised latents. Consequently, the performance after guidance is even worse than without guidance. Following the methodology of Yu et al. (2023), we also attempted to directly predict the corresponding ${\bm{x}}_{0}$ from the intermediate ${\bm{x}}_{t}$ and then use $r({\bm{x}}_{0},y)$ to backpropagate gradients. However, we found that this method cannot generate normal videos.
Table 7: Reward guidance using only MQ rewards on TA-Hard. The reward model trained with noised latents guides the generation effectively, while the model trained on cleaned latents fails to provide meaningful gradient guidance for noised latents.
VQ MQ TA
VDM w/o noise 37.1 38.6 52.9
VDM w/ noise 66.2 74.6 32.4
@@ 6 Conclusion
We have constructed a large-scale preference dataset with 200k human preference annotations, encompassing visual quality, motion quality, and text alignment for modern video generation models. Building on this dataset, we develop VideoReward, a sophisticated multi-dimensional video reward model, and establish the VideoGen-RewardBench benchmark, facilitating more precise and fair evaluation of video generation models. From a unified reinforcement learning perspective, we introduce three alignment algorithms for flow-based video generation models, demonstrating their effectiveness.
Limitations & Future Work
@@
In our experiments, excessive training with Flow-DPO led to a significant deterioration in model quality, despite improvements in alignment across specific dimensions. To prevent this decline, we employed LoRA training. Future work can explore the simultaneous use of high-quality data for supervised learning during DPO training, aiming to preserve video quality while enhancing alignment. Additionally, our algorithms have been validated on text-to-video tasks; future work can extend the validation to other conditional generation tasks, such as image-to-video generation. Moreover, our VideoReward model is vulnerable to reward hacking, wherein human assessments indicate a marked decrease in video quality, yet the reward model continues to assign high scores. This issue arises because the reward function is differentiable, making it susceptible to manipulation. Future research should focus on developing more robust reward models, potentially by incorporating uncertainty estimates and increasing data augmentation. Additionally, there is potential to apply more RLHF algorithms, such as PPO, to flow-based video generation tasks.
Appendix A Details of the Derivation
A.1 Relation beween Velocity and Noise
Lemma A.1.

### v2

@@ 5 Experiments
VideoScore [19] 49.03 71.69 41.80 50.22 47.41 47.72 59.05 51.09 37.24 50.34
LiFT [73] 37.06 58.39 39.08 57.26 47.53 55.97 59.04 54.91 33.79 55.43
VisionRewrd [76] 51.56 72.41 56.77 67.59 47.43 59.03 59.03 60.98 46.56 61.15
@@ 5.1 Reward Learning
We use Qwen2-VL-2B [72] as the backbone of our reward model, trained with BTT loss. Several observations were made during training. First, higher video resolution and more frames generally improved the reward model’s performance. Second, using a stable sampling interval instead of a fixed frame number significantly enhanced motion quality evaluations, especially for videos of varying lengths. In practice, we sample videos at 2 fps, with a resolution of approximately $448\times 448$ pixels while preserving the original asoect ratio. Hyperparameters are in Appendix I.
@@
We evaluate our reward model on two benchmarks targeting different generations of T2V models: (1) VideoGen-RewardBench: Built upon the third-party prompt-video dataset VideoGen-Eval [84], this benchmark targets modern T2V models. We address the lack of human annotations in VideoGen-Eval by manually constructing 26.5k triplets and hiring annotators to provide pairwise preference labels. Annotators also assess overall video quality, serving as a universal label across all dimensions. (2) GenAI-Bench [29]: GenAI-Bench features short (2-seconds) videos generated by pre-Sora-era T2V models, enabling evaluation on earlier-generation outputs. A detailed comparison between the two benchmarks is provided in Appendix H.1. We evaluate our reward model against existing baselines, including VideoScore [19], as well as two concurrent works: LiFT [73] and VisionReward [76]. Consistent with practices in LLM evaluation [35], we use pairwise accuracy, reporting both ties-included [11] and ties-excluded accuracy. We calculate overall accuracy on GenAI-Bench and dimension-specific (VQ, MQ, TA) accuracy on VideoGen-RewardBench. Additional evaluation details are in Appendix H.2 and Appendix H.3
Main Results.
Tab. 2 presents the pairwise accuracy across both benchmarks. VideoScore performs well on GenAI-Bench but fails on VideoGen-RewardBench, indicating poor generalization to modern T2V models. LiFT improves over VideoScore on modern videos but remains below 60% accuracy, showing limited pairwise discrimination ability. VisionReward demonstrates competitive performance on GenAI-Bench but underperforms on VideoGen-RewardBench, especially on visual and motion dimensions under ties-included settings. This drop stems from its difficulty in assessing the improved fidelity and motion smoothness of modern outputs. In contrast, our method VideoReward outperforms all other models on VideoGen-RewardBench, showcasing its strong alignment with human preferences on modern T2V generations. Moreover, despite being trained on a disjoint dataset (see Fig. 10), it still achieves comparable performance on GenAI-Bench, indicating robust generalization across different eras of T2V models. Ablation studies are provided in Appendix E.
5.2 Video Alignment
Training Setting.
Our pretrained model $p_{\text{ref}}$ is an internal, research-purpose video generation model based on Transformer architecture [53], which is trained using rectified flow (see Appendix 7 for details). Following SD3 [15], all alignment experiments fine-tune the Transformer using LoRA [24]. For training-based alignment methods, including SFT, Flow-DPO, and Flow-RWR, we adopt VideoReward to provide the reward signals. For reward guidance, we employ the latent reward model to generate rewards. For supervised fine-tuning (SFT), we utilize only the “chosen data”. Following Rafailov et al. [61], we employ VideoReward as the ground-truth reward model to simulate human feedback and relabel our training dataset, ensuring that models optimised on these synthetic labels can be evaluated fairly by the same reward function. Hyperparameter settings are provided in the Appendix I.
@@ 5.2 Video Alignment
Figure 4: Comparison of videos generated by the original model and the Flow-DPO aligned model.
Evaluation.
We evaluate performance using both automatic and human assessments. For automatic evaluation, we measure the win rate (via VideoReward) and the Vbench score. The win rate is the proportion of cases where VideoReward assigns a higher reward to the aligned model than to the pretrained model. Vbench is a fine-grained T2V benchmark that assesses Quality and Semantic alignment. For human evaluation, each sample is reviewed by two annotators, with a third annotator resolving disagreements. Identical random seeds are used across methods for strict comparability. Our evaluation uses prompts from Vbench, VideoGen-Eval, and a new TA-Hard set that stresses complex semantics, since we notice that the Vbench and VideoGen-Eval prompts are relatively straightforward in terms of text alignment. Appendix L lists a subset of TA-Hard prompts.
Figure 5: Human evaluation of Flow-DPO aligned model vs. pretrained model on VideoGen-Eval.
Figure 6: Accuracy of time-dependent $\beta_{t}$ v.s. constant $\beta$ for TA.
Multi-dimensional Alignment.
We use linear scalarization [40], $r=\frac{1}{3}(r_{\text{vq}}+r_{\text{mq}}+r_{\text{ta}})$, to aggregate multi-dimensional preferences into a single score, and forming a dataset $D=\{({\bm{y}},{\bm{x}}^{w},{\bm{x}}^{l})\}$ where $r({\bm{x}}^{w},{\bm{y}})>r({\bm{x}}^{l},{\bm{y}})$. Table 3 shows that Flow-DPO with a constant $\beta$ significantly improves over the pretrained model across all three dimensions and outperforms both SFT and Flow-RWR. In contrast, Flow-DPO with timestep-dependent $\beta$ underperforms the pretrained model on TA, despite high VQ/MQ reward win rates due to reward hacking. Meanwhile, the constant-$\beta$ variant achieves high VQ/MQ scores without such hacking issues, suggesting it learns TA more effectively. We discuss this further in Sec. 5.2. Figures 6 confirms these findings in human studies.
@@
We also investigate the ability of different methods on a specific task: TA. Tab. 4 shows that Flow-DPO with a constant $\beta$ achieves best performance across all datasets.
Reward Guidance.
We apply linear scalarization to combine dimension rewards into a weighted sum and backpropagate gradients to the noised latent with guidance strength $w=100$. Table 5 shows that users can steer generation toward custom trade-offs by simply choosing custom weights.
Table 5: Reward guidance on VideoGen-Eval.
@@
We meticulously adapted diffusion-DPO to flow-based models, resulting in Equation 4.2, where $\beta_{t}=\beta(1-t)^{2}$. Figure 6 shows that under various $\beta$ values, a constant $\beta$ consistently outperforms the timestep-dependent variant. This is likely because varying $\beta$ across timesteps leads to uneven training [18], given that T2V models use shared weights for different noise levels.
@@ 6 Conclusion
We present a large-scale preference dataset of 182k human annotations covering visual quality, motion quality, and text alignment for modern video generation models. Building on this dataset, we introduce VideoReward, a multi-dimensional video reward model, and establish the VideoGen-RewardBench benchmark for more accurate and fair evaluations. From a unified reinforcement learning perspective, we further propose three alignment algorithms tailored to flow-based video generation, demonstrating their effectiveness in practice.
7 Limitations & Future Work
@@
In our experiments, excessive training with Flow-DPO led to a significant deterioration in model quality, despite improvements in alignment across specific dimensions. To prevent this decline, we employed LoRA training. Future work can explore the simultaneous use of high-quality data for supervised learning during DPO training, aiming to preserve video quality while enhancing alignment. Additionally, our algorithms have been validated on text-to-video tasks; future work can extend the validation to other conditional generation tasks, such as image-to-video generation. Moreover, despite our efforts like incorporating tie annotations into the modeling, our VideoReward model still carries a potential risk of reward hacking, where human assessments indicate a marked decrease in video quality, yet the reward model continues to assign high scores. This issue arises because the reward function is differentiable, making it susceptible to manipulation. Future research should focus on developing more robust reward models, potentially by incorporating uncertainty estimates and increasing data augmentation. Additionally, there is potential to apply more RLHF algorithms, such as PPO, to flow-based video generation tasks.
Acknowledgements
This work was supported by the JC STEM Lab of AI for Science and Engineering, which is funded by The Hong Kong Jockey Club Charities Trust and the Research Grants Council of Hong Kong (Project No. CUHK14213224). It was also supported by the National Natural Science Foundation of China (Grant No. 62576191).
Appendix A Related Works
Evaluation and Reward Models.
Evaluation models and reward models play a pivotal role in aligning generative models with human preferences. Earlier approaches and benchmarks Huang et al. [2023], Liu et al. [2024c], Huang et al. [2024] relied on metrics like FID Heusel et al. [2017] and CLIP scores Radford et al. [2021] to assess visual quality and semantic consistency. Recent works Wu et al. [2023], Xu et al. [2024b], Kirstain et al. [2023], Zhang et al. [2024b], Liang et al. [2024a] have shifted towards utilizing human preference datasets to train CLIP-based models, enabling them to predict preference scores with improved accuracy. With the advent of large vision-language models (VLMs) Achiam et al. [2023], Wang et al. [2024a], their powerful capabilities in visual understanding and text-visual alignment make them a natural proxy for reward modeling. A common approach involves replacing the token classification head of VLMs with a regression head that predicts multi-dimensional scores for diverse evaluation tasks.
Two main learning paradigms have emerged based on the type of human annotation. The first paradigm relies on point-wise regression, where the model learns to fit annotated scores He et al. [2024] or labels Xu et al. [2024a] directly. Another paradigm focuses on pair-wise comparisons Wu et al. [2025], leveraging Bradley-Terry (BT) Bradley and Terry [1952] loss or rank loss to model relative preferences, which is largely unexplored for video reward model. Beyond these methods, some works Wang et al. [2024b] also leverage the intrinsic reasoning capabilities of VLMs through VLM-as-a-judge Li et al. [2023], Lin et al. [2024b], where VLMs are adopted to generate preference judgments or scores in textual format through instruction tuning. Despite these promising advances, most existing video reward models primarily focus on legacy video generation models, typically from the pre-Sora OpenAI [2024] era, which are constrained by short video durations and relatively low quality. Furthermore, the technical choices underlying the vision reward models remain underexplored. Our work seeks to address these limitations by focusing more on modern video generation models and investigating a broader range of reward modeling strategies.
Alignment for Image & Video Generation.
In large language models, Reinforcement Learning from Human Feedback (RLHF) improves alignment with human preferences, enhancing response quality Ouyang et al. [2022], Jaech et al. [2024]. Similar methods have been applied to image generation, using reward models or human preference data to align pretrained models. Key approaches include: (1) direct backpropagation with reward signals Prabhudesai et al. [2023], Clark et al. [2023], Xu et al. [2024b], Prabhudesai et al. [2024]; (2) Reward-Weighted Regression (RWR) Peng et al. [2019], Lee et al. [2023], Furuta et al. [2024]; (3) DPO-based policy optimization Rafailov et al. [2024], Wallace et al. [2024], Liang et al. [2024b], Dong et al. [2023], Yang et al. [2024a], Liang et al. [2024b], Yuan et al. [2024b], Liu et al. [2024b], Zhang et al. [2024a], Furuta et al. [2024]; (4) PPO-based policy gradients Schulman et al. [2017] Black et al. [2023], Fan et al. [2024], Zhang et al. [2024c], Chen et al. [2024a]; and (5) training-free alignment Yeh et al. [2024], Tang et al. [2024], Song et al. [2023]. These methods have successfully aligned image generation models with human preferences, improving aesthetics and semantic consistency. They focus on improving the accuracy of rewards on noised images using reward models trained on clean images, whereas our Flow-NRG directly trains noise-aware reward models to obtain accurate gradients. Our work applies the DPO algorithm Rafailov et al. [2024], Wallace et al. [2024] to flow matching in video generation. Concurrent work Domingo-Enrich et al. [2024] also explores similar things in image generation. However, they reports worse performance for the DPO-aligned model compared to the unaligned one. We argue that the originally derived Flow-DPO algorithm imposes a stronger KL constraint at lower noise steps, resulting in suboptimal performance. In contrast, using a constant KL constraint significantly improves performance on certain tasks. Some prior work explores aligning video generation models using direct backpropagation with differentiable rewards Yuan et al. [2024a], Li et al. [2024b, a], Prabhudesai et al. [2024], often relying on image reward models Kirstain et al. [2023], Wu et al. [2023]. However, these approaches cannot be directly applied to modern T2V generation with a VLM-based video reward and large VAE decoders, as they exceed the memory limits of existing GPU setups, requiring specialized engineering techniques to handle.
Discussion with Concurrent Video Alignment Works.
Several concurrent works also explore aligning video generation models using feedback.  Furuta et al. [2024] derives a unified probabilistic objective for offline RL fine-tuning of text-to-video models for DPO and RWR. VideoDPO Liu et al. [2024b] introduces a re-weighting factor on the Diffusion-DPO loss to adjust the impact of each pair, encouraging the model to learn more effectively from pairs with clearer distinctions. VisionReward Xu et al. [2024a] ensures that all dimensions of the chosen data outperform those of the rejected data, addressing the issue of confounding factors in preference data. OnlineVPO Zhang et al. [2024a] presents an online DPO algorithm to tackle off-policy optimization. LIFT Wang et al. [2024b] proposes applying RWR on synthesized datasets while simultaneously performing supervised learning on real video-text datasets, as synthesized videos often suffer from low temporal consistency. All of these works use Diffusion-DPO or Diffusion-RWR and focus on aligning diffusion-based video generation models, where the videos in the preference datasets are either generated by earlier open-source models or use image reward models directly. In contrast, our work explores alignment techniques for advanced flow-based video generation. We extend Diffusion-DPO into Flow-DPO, but our derivation reveals that the parameter $\beta$ (the KL divergence constraint) in Flow-DPO is timestep-dependent, which leads to suboptimal performance on certain tasks. However, fixing $\beta$ resolves this issue. SPO Liang et al. [2024b] also assumes that the preference order between pairs of images remains consistent when adding the same noise, and constructs win-lose pairs for noised images, proposing step-aware preference optimization. Our work differs from SPO in that while SPO focuses on improving the training method DPO, our Flow-NRG specifically targets training noise-ware reward model on noised videos.
Appendix B Details of the Derivation
B.1 Relation beween Velocity and Noise
Lemma B.1.